OpenStudy (anonymous):

Determine the fourth roots of 16 in the form x+iy, where x and y are not trigonomentric functions. It counts 18 marks. Where do I begin?

OpenStudy (anonymous):

make a circle

OpenStudy (anonymous):



OpenStudy (anonymous):

since \(2^4=16\) you have one root right away, namely \(2\) then divide the circle up in to 4 equal parts, one of them being at \(2\)

OpenStudy (anonymous):



OpenStudy (anonymous):

those are your roots, of course we could do it with algebra as well\[x^4=16\\ x^4-16=0\\ (x^2+4)(x^2-4)=0\\ (x^2+4)(x+2)(x-2)=0\]

OpenStudy (anonymous):

Wow. Thank you. And that really counts 18 marks? :)

OpenStudy (anonymous):

i don't know, it is not my question hope the answers are clear